# Modeling Extensive Defects in Metals through Active Machine Learning and Automated Configuration Reconstruction

### v1

@@ RESULTS
We further evaluate the uncertainty quantification of our defect pool during a continuous nanoindentation simulation, illustrated in Figs 3(g-i). Figs 3(g, h) display the distribution of defect atoms color-coded by $\gamma$ values at the maximum depth (4 nm), as calculated using B and B+C, respectively. For the case of B, several regions exhibit $\gamma$ values above 1, particularly in regions involving dislocation interactions and surface regions, as indicated by red arrows in Fig. 3(g). In contrast, all defect atoms analyzed with B+C demonstrate $\gamma$ values lower than 1 (Fig. 3(h)), indicating significantly lower uncertainty. Additionally, we calculate the maximum $\gamma$ values ($\gamma_{\text{max}}$) among all defect atoms for each snapshot throughout the nanoindentation process. Comparisons of $\gamma_{\text{max}}$ for B, GAP, B+C, and B+P reveal that B and GAP consistently show $\gamma_{\text{max}}$ values above 2, reflecting high uncertainty throughout the nanoindentation process. Conversely, both B+C and B+P maintain $\gamma$ values around 1, with B+C exhibiting slightly lower $\gamma$ values than B+P. This difference is attributed to B+C having a higher proportion of surface atoms, which are prevalent in open-surface nanoindentation simulations. Nonetheless, these results indicate that our defect pool encompasses all the plastic LAEs during nanoindentation, including dislocation nucleation from open surfaces and complex multiplication beneath the indenter.
Figure 3: Uncertainty quantification of various defects in BCC W based on different dataset. a-c Three types of dislocation loop, $<100>$ interstitial loop, $<111>$ glide loop, and $<111>$ climb loop. Left panels are results from B, and right panels are results from B+C. Blue lines are edge dislocation segments, and red lines are screw segments. d Polycrystal model. e, f The probability density function (PDF) of extrapolation grade ($\gamma$) at 300 K and 2000 K. g, h Defect structures at the maximum indentation depth from B and B+C. i Evolution of the maximum extrapolation grade ($\gamma_{\text{max}}$) during the whole indentation process. All atoms in a, b, c, d, g, h are color-coded by $\gamma$. Red arrows in b, c, g indicate the atoms with high uncertainty. A value of $\gamma$ between 0 and 1 indicates interpolation, while a $\gamma$ greater than 1 suggests extrapolation.
To further highlight the greater diversity of LAED in the B+P dataset compared to B+C, we calculate the extrapolation grade $\gamma$ for configurations within the GAP dataset, evaluated by both B+C and B+P. The comparative results are illustrated in Fig. S3a, where we plot the $\gamma$ values for B+C and B+P against each other. Our analysis reveals that the B+P dataset includes 59 configurations with $\gamma$ values exceeding 5, notably, none exceeding 20. In contrast, the B+C dataset shows 74 configurations with $\gamma$ values over 5, including 8 configurations where $\gamma$ surpasses 20. This stark contrast underscores that B+P offers a richer diversity of LAEs than B+C. Interestingly, Fig. S3a also indicates that certain LAEs present in the GAP dataset are missing in B+P. These absent LAEs are primarily configurations of dumbbell interstitials, which are particularly relevant to simulations of radiation damage. Our current EIP-AL procedure in Fig. 1a does not encompass these scenarios, resulting in their absence. Nonetheless, incorporating these specific LAEs into our model through EIP-AL and PCC-GCMC would be straightforward. Furthermore, we calculate the extrapolation grade $\gamma$ for configurations within the dataset B+P based on the dataset GAP, as depicted in Fig. S3b. It is observed that 14.8% of configurations within B+P are absent in GAP. This significant absence explains the high uncertainty observed for the dataset GAP in Fig. 3, indicating that the existing GAP dataset and the resultant MLIP are not adequately equipped for general defect modeling of BCC W. This finding underscores the necessity of our study and the development of a more comprehensive dataset such as B+P, which better addresses the complexities of large-scale simulations.
MLIP refinement via on-the-fly active learning
@@
It should be noted that all LAEDs discussed above are generated using the EAM-Zhou potential  (?) and the MEAM potential  (?). The applicability of the MLIP-0 models developed based on these LAEDs for modeling extended defects in large-scale simulations, particularly within MLIP simulations, remains uncertain. To address this issue, we employ the traditional on-the-fly AL (OTF-AL) implemented in MLIP-3  (?). This process allows us to monitor the extrapolation grade $\gamma$ during MLIP-based simulations and to identify unknown LAEDs. These activities correspond to Step 3 in Fig. 1c.
@@
Initially, we focus on the high-temperature relaxation of a polycrystal. Although Fig. 3e illustrates that our dataset constructed via the initial EIP-AL encompasses all LAEDs in a polycrystal relaxed by the MEAM potential at 300 K, Fig. 3f reveal that these LAEDs do not adequately represent all LAEDs at 2000 K. Consequently, EIP-AL is deemed ineffective for this high temperature. Here, we engage OTF-AL to accommodate polycrystal relaxation at various temperatures: 300 K, 900 K, 1200 K, and 2000 K, starting with the dataset B+P. In the initial AL loop, no new configurations are generated at 300 K and 900 K. However, the simulation halts at 1200 K due to the loss of atoms, prompting the selection of 21 new configurations with $\gamma$ values slightly greater than 2. These clusters are then transformed into periodic configurations via PCC-GCMC and reintegrated into our training database for the development of an updated MLIP. In the subsequent loop, the relaxation simulations proceed smoothly across all the temperatures. During the 2000 K relaxation, an additional 25 new configurations are identified. These activities ensure that we now possess a comprehensive defect pool necessary for modeling general GBs in a random polycrystal, all of which are embedded in periodic configurations via PCC-GCMC. All these configurations are collectively combined as the dataset OTF-AL-GB. To further asses the uncertainty of our MLIP model trained by B+P+OTF-AL-GB, we use ensemble learning to quantify the uncertainty of a polycrysal relaxed at 300 K and 2000 K (Fig. S4). It is seen that the uncertainty is significantly reduced using the new MLIP models. The details of the ensemble learning are provided in Supplementary Note 2.
We next explore a new application: crack propagation in BCC W across four different crack systems. It is important to note that the initial EIP-AL in Fig. 1a did not include fracture simulations. Consequently, we commence with conducting crack propagation simulations using both EAM-Zhou  (?) and MEAM potentials  (?), known for their wide use in fracture analysis. Critical snapshots of crack propagation are presented in Figs 4a-d using EAM and Figs 4e-h using MEAM. MEAM, specifically tailored for better fracture simulation  (?), demonstrates distinct fracture behaviors compared to EAM. For instance, EAM simulations reveal a BCC-to-FCC phase transition at the crack tip within the (010)[001] crack system (Fig. 4a), while MEAM predominantly shows brittle cleavage (Fig. 4e), although a similar phase transition appears in later stages, possibly influenced by boundary conditions. Additionally, EAM captures dislocation emission in the crack systems $(\bar{1}10)[001]$ and $(\bar{1}10)[\bar{1}\bar{1}1]$ (Figs 4b, d), in contrast to the brittle cleavage observed with MEAM in these systems (Figs 4f, h). Both potentials, however, depict deformation twinning at the crack tip in the crack system $(112)[1\bar{1}0]$ (Figs 4c, g). To refine our MLIPs, EIP-AL is applied to simulation outputs from both EAM and MEAM, resulting in the selection of 14 new defect clusters crucial for fracture simulations. These clusters are then processed using PCC-GCMC to create periodic configurations suitable for further DFT simulations, forming the new dataset, EIP-AL-fracture. Building upon this, a new MTP is trained incorporating the datasets B+P+OTF-AL-GB+EIP-AL-fracture. Subsequent OTF-AL for fracture simulations across the four crack systems leads to the selection of 25 additional configurations. After converting these into periodic configurations via PCC-GCMC and calculating them with DFT, they are integrated into the comprehensive datasets, culminating in B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture.
To demonstrate the reduction in uncertainty achieved by integrating the EIP-AL-fracture and OTF-AL-fracture datasets, we apply ensemble learning to calculate the standard deviations of atomic forces for critical snapshots during the OTF-AL of fracture simulations across four crack systems. These calculations are conducted using five independent MTPs developed from datasets B+P+OTF-AL-GB and B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture, respectively. The results are displayed in Figs 4i-p. The analysis reveals that, prior to including crack-related LAEDs, there is significant uncertainty at the surfaces of crack tips and in dislocations nucleated from these crack tips. However, after incorporating the EIP-AL-fracture and OTF-AL-fracture datasets, the uncertainty for all atoms is markedly reduced. This improvement is showcased by lower standard deviations in atomic forces, highlighting the effectiveness of the combined EIP-AL and PCC-GCMC approaches in adapting to new applications such as fracture simulations. These findings underscore that the final datasets, enriched with a comprehensive defect genome, offer robust capabilities for simulating general plastic deformations and crack propagation in BCC W, providing a solid foundation for accurate modeling across various loading configurations and boundary conditions.
@@
Finally, we consolidate all datasets in Table S1 and train a series of MTPs at various levels, alongside a complex ACE potential. Given the high accuracy of ACE in training performance and its predictive capability for fundamental properties of BCC W, coupled with its computational efficiency (Table S2), we select ACE as the definitive MLIP model for conducting subsequent large-scale simulations.
Figure 4: EIP-AL and OTF-AL of fracture simulations. a-h Simulation results of four crack systems using EAM and MEAM potentials. i-l Uncertainty quantification of defects near the crack tip based on MLIPs trained with B+P+OTF-AL-GB. m-p Uncertainty quantification of defects near the crack tip based on MLIPs trained with B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture. All atoms in a-h are color-coded by common neighbor analysis: blue atoms have a BCC lattice, green atoms have an FCC lattice, and white atoms have an unrecognized lattice. All atoms in i-p are color-coded by the standard deviation of atomic force.
MLIP applications in nanoindentation and bicrystal compression
@@
We initiate the application of our ACE potential by simulating nanoindentation in single crystal W. Although nanoindentation has been extensively studied using EAM potentials  (?, ?, ?), there are relatively few simulations employing MLIPs. This scarcity is likely due to the challenges associated with accounting for dislocation nucleation from open surfaces and handling complex dislocation interactions, which are difficult to simulate directly with DFT. In a recent study, a tabulated GAP potential was used for nanoindentation of single crystal W  (?), but the uncertainty associated with these simulations was not detailed. Meanwhile, an advanced neural network potential was specifically developed for nanoindentation in single crystal Mo  (?). In Figs 3(g-i), we demonstrate that the LAEDs obtained by the initial EIP-AL successfully encompass all the plastic events observed during nanoindentation simulations conducted with EAM. Building on this foundation, we now present the simulation results obtained with our final ACE potential and examine the uncertainty.
@@
The simulation setup is aligned with the parameters described in Ref  (?), depicted in Fig. 5a. The indentation is conducted along the [100] direction, which is free from boundary constraints, while the other two axes maintain periodic boundary conditions. The total system spans dimensions of 38 $\times$ 38 $\times$ 32 $\text{nm}^{3}$, encompassing 2,832,200 atoms. To stabilize the substrate, the bottom layers, with a thickness of 6.4 Å, are fixed. Additionally, several layers, totaling a thickness of 2.56 nm, function as thermostat layers to dissipate heat generated during the nanoindentation process. The remaining atoms are designated as Newtonian layers, where the indentation process occurs. The indentation is performed at a velocity of 20 m/s, using a spherical indenter with a radius of 6 nm. Fig. 5b displays the force-depth curve, reaching a maximum depth of 3.8 nm. The corresponding microstructures at this depth are depicted in Fig. 5c, where all defect atoms are color-coded based on the extrapolation grade ($\gamma$). Notably, all defects exhibit $\gamma$ values lower than 1.2, affirming the low uncertainty in modeling the complex dislocation network generated by the ACE potential. The dislocation pattern reveals prismatic loops advancing along 111 slip directions. This pattern, which includes the process of shear loop formation and cross-slip repeating beneath the indenter tip, aligns with results from previous studies using tabGAP and experimental observations  (?). Additionally, the slip traces and pileup, detailed in Fig. 5d, resonate with findings from the customized neural network potential employed for Mo nanoindentation  (?). These results not only validate the effectiveness of the ACE in capturing intricate deformation mechanisms but also demonstrate its consistency with established models and experimental data.
@@
Our simulations reveal a new mechanism for the onset of plasticity during nanoindentation, as analyzed using BCC Defect Analysis (BDA)  (?). The outcomes of this analysis are illustrated in Fig. 5e, where nano-screw dislocations, deformation twinning, 110 planar faults, surface features, and vacancies are identified. Specifically, the sequence of events leading to the initiation and evolution of plasticity is marked by five critical points A-E, labeled in Fig. 5b and detailed in Fig. 5e. At point A, plasticity begins with the formation of a local hexagonal close-packed (HCP) region. The first yield, at point B, is associated with the formation of three intersecting twin boundaries (TBs). At point C, the vertical TB extends deeper into the material while the left one disappears. At point D, the vertical TB further advances, and the right TB transitions into a regular GB. Finally, at point E, the extension of the vertical TB halts. These results indicate that while deformation twinning is critical at the initial stages of nanoindentation (from B to D), dislocation nucleation and multiplication subsequently take precedence, governing the later stages of deformation (from D to E).
Figure 5: Simulation results of nanoindentation in BCC W. a Simulation setup. b Load-depth curves with five critical points A-E labeled. c The defect pattern at maximum depth, with all atoms color-coded by the extrapolation grade $\gamma$. Arrows indicate the {111} slip directions. d Positions of atoms on the top surface; the slip trace is indicated by a circle. e BCC defect analysis (DBA) at five critical points A-E.
We further apply our ACE potential to investigate dislocation-GB interactions within two distinct spherical bicrystals, each characterized by either a low-angle (GB1) or a high-angle twist GB (GB2). The specific lattice orientations for these bicrystals are illustrated in Fig. 6a. Both bicrystal configurations comprise 258,043 atoms. To facilitate dislocation nucleation and subsequent interactions with the GBs, a rigid plane is displaced downward, compressing the bicrystal. Following each loading increment of 0.25 Å, the system is allowed to relax for 5 ps, achieving a loading velocity of 5 m/s. These simulations are conducted using the NVT ensemble at a temperature of 5 K, ensuring precise control over the thermal conditions and mechanical deformations. Fig. 6b displays the atomic structure of the two distinct GBs. It is observed that the low-angle twist GB is characterized by a dislocation network. The mechanical responses of the two bicrystals, labeled BC1 and BC2, under compression are meticulously documented through various metrics in Figs S5 and Fig. 6c. These figures display the contact area-displacement curves, force-displacement curves, and stress-displacement curves, respectively. Analysis of these curves reveals four distinct behavioral regimes, S1 through S4, each demonstrating unique characteristics between the two bicrystals. In particular, in regimes S1 and S2, BC1 and BC2 exhibit comparable contact areas, as shown in Fig. S5a. However, in regimes S3 and S4, BC2 presents a larger contact area. Concurrently, Fig. S5b illustrates that in regime S1, the forces exerted by BC1 and BC2 are equivalent, whereas in regimes S2 and S4, BC2 exerts greater force than BC1, and conversely, in regime S3, BC1 surpasses BC2 in force. The stress, defined as the ratio of force to contact area, is depicted in Fig. 6c. Here, BC1 and BC2 display comparable stress in regimes S1 and S4. BC2 exhibits greater stress in regime S2, while BC1 shows higher stress than BC2 in regime S3. Furthermore, we have computed the maximum extrapolation grade ($\gamma_{\text{max}}$) across the entire compression process, detailed in Fig. S5c. This assessment confirms the robustness of our ACE model in accurately predicting the forces and energy throughout bicrystal deformation, even under extreme strain conditions.
We conduct atomic shear strain analysis and extract dislocation lines to understand the distinct deformation mechanisms in the two bicrystals, BC1 and BC2. The detailed analyses of atomic slip histories for BC1 and BC2 at four critical points A-D, as labeled in Fig. 6c, are presented in Figs 6d ,e. Fig. S6 displays the extracted dislocation lines within each bicrystal. In the initial stage, regime S1, both bicrystals exhibit overlapping curves in Fig. 6c and Figs S5a, b due to similar behaviors characterized by elastic deformation and dislocation nucleation, which are independent of GB type. In regime S2, BC2 experiences higher forces and stress compared to BC1. This difference is attributed to the ease of dislocation transmission across GB1, altering its dislocation network, as shown in Fig. 6d-A and Fig. S6a-A. Conversely, all dislocations are impeded by GB2, as depicted in Fig. 6e-A and Fig. S6b-A, demonstrating the influence of GB strengthening or the Hall-Petch effect, which contributes to the increased force and stress in BC2. In regime S3, BC1 exhibits higher force and stress than BC2. This is due to the smaller contact area in BC1 compared to BC2 (Fig. S5a). In BC1, deformation is uniformly distributed across the bicrystal (Fig. 6d-B and Fig. S6a-B), whereas in BC2, deformation concentrates in the upper grain (Fig. 6e-B and Fig. S6b-B), leading to a larger contact area and subsequently lower nucleation stress for dislocations in BC2. In regime S4, the stresses of the two bicrystals become comparable. This results from a balance between GB strengthening and nucleation weakening in regime S4 in BC2, as illustrated in Fig. 6d-C and Fig. S6-C. In fact, BC1 behaves similarly to a single crystal throughout the compression process because the low-angle twist GB exhibits a negligible blocking effect on dislocation movement (Fig. 6d-C and Fig. S6a-C). Conversely, in BC2, the high-angle twist GB acts as a robust barrier, significantly impeding dislocation mobility. Even at the end of the compression process (point D), this effect is evident as the upper grain of BC2 is compressed into a flat layer, while the lower grain retains its original spherical shape, as depicted in Fig. 6d-D and Fig. S6b-D. Moreover, significant GB sliding is observed in BC2, particularly highlighted in Figs 6c-C, D. These findings underscore the capability of our ACE model to accurately simulate and analyze complex plastic deformations involving intricate interplays between dislocation motion and GB deformation.
Figure 6: Simulation results of bicrystal compression. a Simulation setup and lattice orientations of upper and lower grains. b GB structures. c Stress-displacement curves of two bicrystals. Four regimes S1-S4 and four critical points A-D are identified. Atomic slip history of the bicrystal d containing a low-angle GB and e containing a high-angle GB. All atoms in d and e are color-coded by the atomic shear strain.

### v2

@@ RESULTS
We further evaluate the uncertainty quantification of our defect datasets during a continuous nanoindentation simulation, illustrated in Figs 3(g-i). Figs 3(g, h) display the distribution of defect atoms color-coded by $\gamma$ values at the maximum depth (4 nm), as calculated using B and B+C, respectively. For the case of B, several regions exhibit $\gamma$ values above 1, particularly in regions involving dislocation interactions and surface regions, as indicated by red arrows in Fig. 3(g). In contrast, all defect atoms analyzed with B+C demonstrate $\gamma$ values lower than 1 (Fig. 3(h)), indicating significantly lower uncertainty. Additionally, we calculate the maximum $\gamma$ values ($\gamma_{\text{max}}$) among all defect atoms for each snapshot throughout the nanoindentation process. Comparisons of $\gamma_{\text{max}}$ for B, GAP, B+C, and B+P reveal that B and GAP consistently show $\gamma_{\text{max}}$ values above 2, reflecting high uncertainty throughout the nanoindentation process. Conversely, both B+C and B+P maintain $\gamma$ values around 1, with B+C exhibiting slightly lower $\gamma$ values than B+P. This difference is attributed to B+C having a higher proportion of surface atoms, which are prevalent in open-surface nanoindentation simulations. These results indicate that our defect pool encompasses all the plastic LAEs during nanoindentation, including dislocation nucleation from open surfaces and complex multiplication beneath the indenter. It should be noted that datasets B+C and B+P exhibit comparable performance in uncertainty quantification analysis in Fig. 3. This is because the key LAEs for these defects are included in the core atoms of both datasets. However, this does not mean that dataset P, constructed using PCC-GCMC, is unnecessary. In the following section, we will show that dataset P outperforms dataset B in predicting unseen LAEs.
Furthermore, we compare the LAEs in the existing GAP dataset with those in B+P and B+C, as detailed in Supplementary Note 1. Our analysis shows that the GAP dataset and its corresponding MLIP are inadequate for effectively modeling general defects in BCC W. Taken together with the results shown in Fig. 3, these insights emphasize the critical role of EIP-GS as an advanced sampling method, crucial for ensuring data diversity and comprehensive coverage..
Direct Validation Through Comparison with DFT Calculations
To further elucidate the superiority of PCC-GCMC and the resultant P-MLIP-0, we undertake additional validation tasks by comparing the energy and forces predicted by our MLIP-0 models against new DFT calculations for various GBs relaxed at temperatures of 300 K, 900 K, and 2000 K, as well as under severe compression conditions. The stable structures of these GBs are referenced from an established study in Ref  (?). The GB structures and comparative results for energy and force between MLIPs and DFT are displayed in Fig. 4. Nine distinct GBs are considered, which transform into FCC lattices, hexagonal close-packed (HCP) lattices, and even an amorphous phase under severe compression, as shown in Fig. 4a. Observations reveal that both C-MLIP-0 (trained by B+C) and P-MLIP-0 (trained by B+P) consistently outperform B-MLIP-0 (trained solely by B) across all validation scenarios. An exception is noted in the energy comparison at 2000 K, where B-MLIP-0 records a lower error than C-MLIP-0 and P-MLIP-0 (6.3 vs 12.2 and 7.2 meV/atom). However, the force RMSE for B-MLIP-0 is significantly higher compared to C-MLIP-0 and P-MLIP-0 (199.3 vs 172.2 and 154.4 meV/Å). These results suggest that incorporating either cluster (C) or periodic (P) datasets enhances the accuracy. Notably, P-MLIP-0 consistently surpasses C-MLIP-0 in both energy and force accuracy, particularly under compression conditions (6.7 vs 10.1 meV/atom for energy; 150.8 vs 173.7 meV/Å for force). This superior performance can be attributed to the inclusion of additional atoms by PCC-GCMC in Fig. 2d, which introduces relevant LAEs absent in B+C, and thus better aligns with the new validation tasks in Fig. 4b. Combining the training performance illustrated in Fig. 2e, we conclude that B+P offers superior training outcomes and higher transferability than B+C, thereby validating its enhanced efficacy in modeling complex defects.
MLIP Refinement
@@
It should be noted that all LAEs discussed above are generated using the EAM-Zhou potential  (?). The applicability of the MLIP-0 models developed based on these LAEs for modeling extended defects in large-scale simulations, particularly within MLIP simulations, remains uncertain. To address this issue, we employ the traditional on-the-fly AL (OTF-AL) implemented in MLIP-3  (?). This process allows us to monitor the extrapolation grade $\gamma$ during MLIP-based simulations and to identify unknown LAEs. These activities correspond to Step 3 in Fig. 1c.
@@
We focus on the high-temperature relaxation of a polycrystal. Fig. 3e shows that our dataset, constructed via the initial EIP-GS, includes all LAEs in a polycrystal relaxed by the MEAM potential at 300 K. However, Fig. 3f reveals that these LAEs do not fully represent those at 2000 K, making EIP-GS ineffective for high temperatures. To address this, we use OTF-AL to model polycrystal relaxation at 300 K, 900 K, 1200 K, and 2000 K, starting with the dataset B+P. In the first AL loop, no new configurations are generated at 300 K and 900 K. However, the simulation at 1200 K stops due to atom loss, leading to the selection of 21 new configurations with $\gamma>2$. These are converted into periodic configurations via PCC-GCMC and added to the training database to develop an updated MLIP. In the next loop, relaxation simulations proceed smoothly across all temperatures, with 25 additional configurations identified at 2000 K. This ensures a comprehensive defect genome for modeling general GBs in random polycrystals, all embedded in periodic configurations via PCC-GCMC. These configurations form the dataset OTF-AL-GB. To assess the uncertainty of our MLIP trained on B+P+OTF-AL-GB, we use ensemble learning (detailed in the Supplementary Note 2) to evaluate a polycrystal relaxed at 300 K and 2000 K (Fig. 5). The results show significantly reduced uncertainty, with most force standard deviations ($\sigma_{z}$) below 0.2 eV/Å, comparable to MLIP force prediction errors. This confirms the high accuracy of the new MLIP in modeling general GBs.
@@
Additionally, we refine our MLIP model by incorporating crack propagation, as detailed in Supplementary Note 3. Our approach effectively captures the local atomic environments (LAEs) associated with various fracture mechanisms. Finally, we consolidate all datasets in Table S1 and train a series of MTPs at multiple levels, alongside a complex ACE potential. Due to ACE’s exceptional training accuracy, predictive capability for fundamental properties of BCC W, and computational efficiency (see the details in the Supplementary Note 4), we designate ACE as the final MLIP model for subsequent large-scale MD simulations.
MLIP applications in nanoindentation
@@
We initiate the application of our ACE potential by simulating nanoindentation in single crystal W. Although nanoindentation has been extensively studied using EAM potentials  (?, ?, ?), there are relatively few simulations employing MLIPs. This scarcity likely arises from the challenges of modeling dislocation nucleation from open surfaces and capturing complex dislocation interactions, which are difficult to simulate directly with DFT. In a recent study, a tabulated GAP potential was used for nanoindentation of single crystal W  (?), but the uncertainty associated with these simulations was not detailed. Meanwhile, an advanced neural network potential was specifically developed for nanoindentation in single crystal Mo  (?). We now present the simulation results obtained with our final ACE potential and examine the uncertainty.
@@
The simulation setup is aligned with Ref  (?), depicted in Fig. 6a. The simulation details are provided in Methods. Fig. 6b displays the force-depth curve, reaching a maximum depth of 3.8 nm. The corresponding microstructures at this depth are depicted in Fig. 6c, where all defect atoms are color-coded based on the extrapolation grade ($\gamma$). Notably, all defects exhibit $\gamma$ values lower than 1.2, affirming the low uncertainty in modeling the complex dislocation network generated by the ACE potential. The dislocation pattern reveals prismatic loops advancing along $\{111\}$ slip directions. This pattern, which includes the process of shear loop formation and cross-slip repeating beneath the indenter tip, aligns with results from previous studies using tabGAP and experimental observations  (?). Additionally, the slip traces and pileup, detailed in Fig. 6d, resonate with findings from the customized neural network potential employed for Mo nanoindentation  (?). These results not only validate the effectiveness of the ACE in capturing intricate deformation mechanisms but also demonstrate its consistency with established models and experimental data.
@@
Our simulations reveal a new mechanism for the onset of plasticity during nanoindentation, as analyzed using BCC Defect Analysis (BDA)  (?). The outcomes of this analysis are illustrated in Fig. 6e, where nano-screw dislocations, deformation twinning, $\{110\}$ planar faults, surface features, and vacancies are identified. Specifically, the sequence of events leading to the initiation and evolution of plasticity is marked by five critical points A-E, labeled in Fig. 6b and detailed in Fig. 6e. At point A, plasticity begins with the formation of a local HCP region. The first yield, at point B, is associated with the formation of three intersecting twin boundaries (TBs). At point C, the vertical TB extends deeper into the material while the left one disappears. At point D, the vertical TB further advances, and the right TB transitions into a regular GB. Finally, at point E, the extension of the vertical TB halts. These results indicate that while deformation twinning is critical at the initial stages of nanoindentation (from B to D), dislocation nucleation and multiplication subsequently take precedence, governing the later stages of deformation (from D to E).